1. A boy with a mass of 110.41 kg on the top of a platform 10 m high preparing for a dive into the pool below. What will be his Kinetic Energy in Joules as he hits the water?
Answer #1

m =110.41 kg , g =9.8 m/s^2 , h =10 m

From conservation of energy

K1+U1 =K2+U2

K1 +0 = 0+mgh = 110.41*9.8*10

K = 1082 J